OpenStudy (anonymous):

which of the following is not a perfect square? A. 36 B. 45 C. 64 D. 81

OpenStudy (anonymous):

45

OpenStudy (anonymous):

Perfect square means it is a square of an integer. For example, 36 is a perfect square because 6 * 6 or 6^2 = 36. Same with 64. 8*8. Find one that is not a square of an integer.

OpenStudy (anonymous):

36 = 6*6 64 = 8*8 81 = 9*9